OpenStudy (anonymous):

if logb(3)=.23 and logb(8)=.71, find logb24

jimthompson5910 (jim_thompson5910):

Hint: log(24) = log(3*8) = log(3)+log(8) for any base

OpenStudy (anonymous):

i'm supposed to round the answer to the nearest thousandth. how do i get that from this?

jimthompson5910 (jim_thompson5910):

log(3) is what?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

they give it as 0.23

jimthompson5910 (jim_thompson5910):

I'm not including the base, but I'm implying log(3) means "log base b of 3"

OpenStudy (anonymous):

i don't know

jimthompson5910 (jim_thompson5910):

log(24) = log(3)+log(8) log(24) = 0.23 + 0.71 <<--- these are given at the top log(24) = 0.94

OpenStudy (anonymous):

than...i really need to review the formulas, because i really do not understand algorithms.

OpenStudy (anonymous):

thanks...

jimthompson5910 (jim_thompson5910):

The idea I was using at the top was that log(xy) = log(x)+log(y)

jimthompson5910 (jim_thompson5910):

you're welcome